Can you help with this story problem?
Mike picked three-fourths of the pears on his father's tree. After Mike finished, Sylvia picked two-thirds of the pears that were left on the tree. After Sylvia finished, there were twenty pears left on the tree. How many pears were on the tree before Mike and Sylvia started picking?
Thank you for your time.

Found 2 solutions by checkley71, ankor@dixie-net.com:
Answer by checkley71(8403)

x=(3/4)x+2/3(1/4)x+25 or x=(3/4)x+(2/12)x+25 or x=(11/12)x+25 or
x-(11/12)x=25 or (1/12)x=25 or x=25*12 or x=300 pears were on the tree when they started picking.
proof Mike picked 3/4*300=225 which left 300-225=75 left for Sylvia.
Sylvia picked 2/3*75=50 which left 75-50=25 apples unpicked.

Answer by ankor@dixie-net.com(22740)

Mike picked three-fourths of the pears on his father's tree. After Mike finished, Sylvia picked two-thirds of the pears that were left on the tree. After Sylvia finished, there were twenty pears left on the tree. How many pears were on the tree before Mike and Sylvia started picking?
:
Let x = number of pears originally on the tree
:
"Mike picked three-fourths of the pears "
Mike picked: (3/4)x, then (1/4)x is left
:
"Sylvia picked two-thirds of the pears that were left"
If Sylvia picked(2/3) then 1/3 are left
:
(1/3)(1/4)x = 20
(1/12)x = 20
x = 20 * 12
x = 240 pears originally on the tree
:
Check: 240 - 3/4(240) = 60
60 - 2/3(60) = 20